OpenStudy (anonymous):

What is the slope of the line? y=-3x+8

OpenStudy (anonymous):

y=mx+b where m represents the slope :)

OpenStudy (anonymous):

So whats that mean? what do I have to do to solve it?

OpenStudy (anonymous):

the number in place of the m is your answer.

OpenStudy (anonymous):

k, how do i get the number?

OpenStudy (anonymous):

y=3x+8 y=mx+b

OpenStudy (anonymous):

ooh is it 3 lol

OpenStudy (anonymous):

y3a.

OpenStudy (anonymous):

m = -3, to be precisely.

OpenStudy (anonymous):

-3 why?

OpenStudy (anonymous):

ooh nvm

OpenStudy (anonymous):

ty